Suppose that W is inversely proportional to V. If W=20 when V=12, find V=15
Answer by Cromlix(4381)

W is inversely proportional to V.
W = k/V
k = constant of proportionality
To find 'k'
W = k/V
20 = k/12
k = 240
Formula:
W = 240/V
W = 240/15
W = 16.
Hope this helps.
:-)